Enjin Vroom Sdn Bhd operates a factory at the Senawang Industrial Zone that manufacture micron filter for diesel engine. The data given below are for costs in year 2018 for producing and selling 5000 units of the micron filter. Currently, each micron filters been sold at RM160.

Direct labour               RM 86000

Direct materials         RM 60000

Variable manufacturing overhead    RM ?

Fixed manufacturing overhead    RM 39000

Variable selling & administrative costs    RM 7000

Fixed selling & administrative costs          RM  12000



a) Given that the manufacturing cost per unit is RM110, calculate the VARIABLE manufacturing overhead costs in year 2018.Format : 597000

b)  Using the  that you have calculated in part (a), compute the following:

i. Total variable costs

Format : 792000
ii.  Total variable costs per unitFormat : 808.5
iii. Total costsFormat : 689000

c) Based on the current selling price set by the company, calculate the following:

i. Profit per unit for the sold micron filter.

Format : 79.8
ii.  Contribution margin per unit of the micron filterFormat : 85.9
iii. Break-even sales amountFormat : 667976
d. Due to change of supplier, the direct materials cost has increased by 10%. The company's new lean manufacturing practice has successfully reduced the fixed manufacturing overhead expenses by RM10,000 per year. Calculate the new breakeven units.Format : 737
Answer #1

Units Produced & Sold = 4100

Selling Price = RM 180 / Unit

Direct labor = RM 98000

Direct Material = RM 60000

Fixed Manufacturing Overhead = RM 30000

Variable Selling & Administrative costs = RM 8100

Fixed Selling & Administrative costs = RM 12000

(a) Manufacturing Cost per Unit = RM 110

Total Manufacturing Cost = RM 110 * 4100 = 451000

Variable Manufacturing Overhead = Total Manufacturing Cost - Direct labor - Direct Material - Fixed Manufacturing Overhead

= RM 451000 - RM 98000 - RM 60000 - RM 30000

= RM 263000

(b) (i) Total Variable Cost = Direct labor + Direct Material + Variable Manufacturing Overhead + Variable Selling & Administrative costs

= RM 98000 + RM 60000 +RM 263000 + RM 8100 = RM 429100

(ii) Variable Cost per Unit = RM 429100 / 4100 = RM 104.66

(iii) Total Cost = Direct labor + Direct Material + Variable Manufacturing Overhead + Variable Selling & Administrative costs + Fixed Manufacturing Overhead + Fixed Selling & Administrative costs

= RM 98000 + RM 60000 +RM 263000 + RM 30000 + RM 8100 + RM 12000 = RM 471100

(c) (i) Total Sales (4100* RM 180) = RM 738000

Total Cost = RM 471100

Profit = Total Sales - Total Cost

= RM 738000 - RM 471100 = RM 266900

Profit per Unit = RM 266900 / 4100 = RM 65.1

(ii) Total Sales = RM 738000

Total Variable Cost = RM 429100

Contribution = RM 738000 - RM 429100 = RM 308900

Contribution per Unit = RM 308900 / 4100 = RM 75.34

(iii) Break Even Sales = Total Cost = RM 471100

(d)

Direct Material Cost increased by 10%

New Direct Material Cost = (110% of RM 60000) = RM 66000

New Fixed Manufacturing Overhead = RM 20000

Break-Even Units = Total Fixed Cost / New Contribution Margin per Unit

= RM 32000 /  RM 73.88 = 433 Units

Note -

New Total Variable Cost = RM 429100 + RM 6000 = RM 435100

Contribution = RM 738000 - RM 435100 = RM 302900

New Contribution per Unit = Contribution / Units = RM 302900 / 4100 = RM 73.88


answered by: ANURANJAN SARSAM